A commercial jet can fly 450 mph in calm air. Traveling with the jet stream, the plane flew 2400 mi in the same amount of time it would take to fly 1600 mi against the jet stream. Find the rate of the jet stream.
Answer by mananth(16946)

jet speed = 450 mph
wind speed given = x mph

speed against wind 450 - x
speed with wind 450 + x

Distance against wind 1600 miles
Distance with wind 2400 miles

t=d/r time against = time with wind -
1600 / ( 450 - x )= 2400 / ( 450 + x )
1600 ( 450 + x )= 2400 ( 450 -x )
720000 + 1600 x = 1080000 - -2400 x
2400 x 1600 x 1080000 + -720000
4000 x = 360000
/ 4000
x = 90 mph stream speed